OpenStudy (anonymous):

help me 8. Which of the following is the equation of the line passing through points A(−5, 2) and B(0, −8)? (1 point)y = −2x + 8y = 2x + 8y = −2x − 8y = 2x − 8

OpenStudy (solomonzelman):

Didn't you ask this already?

OpenStudy (solomonzelman):

Or maybe someone else did, then my bad!

OpenStudy (anonymous):

yes but you didn't tell me which one it is

OpenStudy (solomonzelman):

I told you what to do though You have to put each equation into y=mx+b form, right?

OpenStudy (solomonzelman):

Write this equation down for me real quick, pliz

OpenStudy (solomonzelman):

equations

OpenStudy (anonymous):

8. Which of the following is the equation of the line passing through points A(−5, 2) and B(0, −8)? (1 point) (a.y = −2x + 8) (b.y = 2x + 8) (c.y = −2x − 8) (d.y = 2x − 8)

OpenStudy (solomonzelman):

You have A(−5, 2) B(0, −8) (b/f comma it is the x and after it is the y- saying it is (x,y) right ?

OpenStudy (anonymous):

yes what is the answer d

OpenStudy (anonymous):

yes what is the answer d

OpenStudy (solomonzelman):

did you plug in the x and y values into each equation? which one of them satisfies both A(−5, 2) & B(0, −8) ?

OpenStudy (anonymous):

no but what is the answer ive been on this question for 2 hours

OpenStudy (solomonzelman):

y = −2x + 8 y = 2x + 8 y = −2x − 8 y = 2x − 8 recalling....

OpenStudy (solomonzelman):

C

OpenStudy (anonymous):

13. Find the value of x so that f(x) is parallel to g(x) (1 point) a. 44, b. 25, c. 42, d. 18